OpenStudy (anonymous):

can anyone help me solve this function....g(3)+f(0) please

OpenStudy (anonymous):

define f(x) and g(x)

OpenStudy (anonymous):

substitute the numbers for the functions so what 's the functions

OpenStudy (anonymous):

that's what i wanna know. you must be given f(x) and g(x) to find f(0) and g(3)

OpenStudy (anonymous):

all you do is substitue that in for f(x) and g(x) what school do you go to if you don't mind me asking

OpenStudy (anonymous):

a university in Canada. the initial question you asked is incomplete. where did you find this problem?

OpenStudy (anonymous):

my teacher gave me a problem to do for practice so i kinda know what to do

OpenStudy (anonymous):

wouldnt it just be 3 because 3+0 is three?

OpenStudy (anonymous):

do you have the functions for what f(x) equals?

OpenStudy (anonymous):

like f(x) = 2x+8 or something like that?

OpenStudy (anonymous):

my function is f(x)=3x-2 g(x)=x-9 g(3)+ f(0) help me please

OpenStudy (anonymous):

f(x)=3x-2 g(x)=x-9 g(3)+ f(0) so lets look at the g(3) so that means any time you see an x in the equation, g(x)=x-9, you replace it with 3

OpenStudy (anonymous):

ok so it's going to be 3-9 right

OpenStudy (anonymous):

yep, but don't forget the write it g(3)=3-9

OpenStudy (anonymous):

wow youre smart what school do you go to

OpenStudy (anonymous):

but ok

OpenStudy (anonymous):

do you wana try the f(0) on your own?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

ok so i got f(0)=3(0)-2

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

g(3)=3-9 = -6 f(0)=3(0)-2 = -2 right?

OpenStudy (anonymous):

now do f(0) + g(3)

OpenStudy (anonymous):

you add right

OpenStudy (anonymous):

right

OpenStudy (anonymous):

f(0) + g(3) is the same as -2 + -6 right? because what we got before

OpenStudy (anonymous):

so its -8

OpenStudy (anonymous):

ok so much i gotta headache

OpenStudy (anonymous):

correct, -8 is the answer

OpenStudy (anonymous):

i have to go now. good job

OpenStudy (anonymous):

ok